# The appearance of the regular black hole with a stable Cauchy horizon

### 07 Dec 2023 v1

@@ V Conclusions and discussion
Many people believe that singularities do not exist within the realm of nature. Therefore, regular black holes which removing the curvature singularity at its center attract a lot of interest. Although they have no singularities, they usually have a unstable Cauchy horizon accompanied by the mass inflation. This instability eventually turns the Cauchy horizon into a null singularity which ensures the validity of SCCC. However, a regular black hole with a stable Cauchy horizon is proposed Carballo-Rubio:2022kad . This inner-extremal regular black hole has a vanishing inner surface gravity, which avoids the mass inflation. Of course, this is in contradiction with the SCCC and violates the predictability of classical theories.
In this paper, we study the global causal structure of this regular black hole, and draw its Penrose diagram. The stable Cauchy horizon allows spacetime to repeat infinitely and makes it possible for the photons to pass through the horizons safely. Thus the photons from a companion universe can fly to our universe through the black hole-white hole channel. This process will result in a new appearance of the regular black hole. Therefore, using a thin accretion disk as the light source, we study the image of this regular black hole by the ray-tracing method. When the photons are emitted and received at the same universe, the image is similar to that of the Schwarzschild black hole. Nevertheless, as for the case the accretion disk located around the companion black hole in the preceding universe and the observer located in our universe, the novel multi-ring structure occurs. There are many extra bright rings in the image, and some of them are distinctly inside the shadow region. Besides, these rings contain the information about the background geometry, so we may obtain the parameters in the metric from the image. For this reason, we depict the images with different parameters of the regular black hole and locations of the accretion disk.
@@
Although the regular black hole we studied has a novel image, a precise physical process that leads to the formation of such black hole remains a mystery. This puts into question whether such a black hole can exist and whether it can be detected in astronomical images. On the other hand, the two-dimensional dilaton black hole and the black hole in higher-derivative gravity theories may have degenerate Cauchy horizons with vanished surface gravities, which have been found by Cai more than twenty years ago Cai:1995nt ; Cai:1998yp . Their Cauchy horizons are stable against the classical perturbations. This provides us with the assurance that we can find a regular black hole with a stable Cauchy horizon in certain theoretical models. This will be the focus of our subsequent research.

### 23 Feb 2024 v2

@@ V Conclusions and discussion
Many people believe that singularities do not exist within the realm of nature. Therefore, regular black holes which removing the curvature singularity at its center attract a lot of interest. Although they have no singularities, they usually have a unstable inner horizon accompanied by the mass inflation. This instability eventually turns the inner horizon into a null singularity which ensures the validity of SCCC. However, a regular black hole with a stable inner horizon is proposed Carballo-Rubio:2022kad . This inner-extremal regular black hole has a vanishing inner surface gravity, which avoids the mass inflation. Of course, this is in contradiction with the SCCC and violates the predictability of classical theories.
In this paper, we study the global causal structure of this regular black hole, and draw its Penrose diagram. The stable inner horizon allows spacetime to repeat infinitely and makes it possible for the photons to pass through the horizons safely. Thus the photons from a companion universe can fly to our universe through the black hole-white hole channel. This process will result in a new appearance of the regular black hole. Therefore, using a thin accretion disk as the light source, we study the image of this regular black hole by the ray-tracing method. When the photons are emitted and received at the same universe, the image is similar to that of the Schwarzschild black hole. Nevertheless, as for the case the accretion disk located around the companion black hole in the preceding universe and the observer located in our universe, the novel multi-ring structure occurs. There are many extra bright rings in the image, and some of them are distinctly inside the shadow region. Besides, these rings contain the information about the background geometry, so we may obtain the parameters in the metric from the image. For this reason, we depict the images with different parameters of the regular black hole and locations of the accretion disk.
@@
Although the regular black hole we studied has a novel image, a precise physical process that leads to the formation of such black hole remains a mystery. This puts into question whether such a black hole can exist and whether it can be detected in astronomical images. On the other hand, Cai has found the two-dimensional dilaton black hole and the black hole in higher-derivative gravity theories which have degenerate Cauchy horizons with vanished surface gravities more than twenty years ago Cai:1995nt ; Cai:1998yp . Consequently, their Cauchy horizons are immune from instability. This provides us with the assurance that we can find a regular black hole with a stable inner horizon in certain theoretical models. Although, their inner horizons are suggested to exhibit instability when taking into account the quantum effects Hollands:2019whz , whether quantum effects significantly alters the behavior of astronomical massive black holes is still an issue of ongoing debate. Therefore, the multi-ring structure could serve as a probe into the potential influence of quantum effects on black hole stability. The emergence of this structure may suggest that quantum effects play a less significant role than previously thought in these massive objects. This will be the focus of our subsequent research.